# We Need to Measure Data Diversity in NLP — Better and Broader

### v1

@@ 1 Introduction
Language models seem to exhibit remarkable natural language understanding and generation capabilities. Yet numerous studies have pointed towards serious flaws: they encode societal biases (Gallegos et al., 2024; Hofmann et al., 2024) and are sensitive to spurious correlations (Du et al., 2023). Many of these weaknesses have been traced back to the training data (Feng et al., 2023; McCoy et al., 2024). Moreover, as the size of these training datasets continues to increase, improvements in task performance show diminishing returns (Tirumala et al., 2023). Data size is often prioritized at the expense of data quality (Paullada et al., 2021). However, there is an increasing effort into analyzing the quality and composition of NLP datasets (Albalak et al., 2024; Elazar et al., 2024; van Noord et al., 2025).
@@
Unfortunately, the concept of “data diversity” is challenging to define and measure, and researchers are rarely explicit about what they mean with data diversity (Zhao et al., 2024). But without clear definitions and measurements, it is difficult to make progress. Previous papers that advocate for measuring data diversity have left important aspects unaddressed. Mitchell et al. (2023) emphasise the importance of measuring data characteristics, though diversity in NLP data was not their main focus. Zhao et al. (2024) highlight the need to measure data diversity; however, their discussion mostly focuses on considerations during dataset construction, rather than diversity measurement itself. Shaib et al. (2025) call for standardizing text diversity measurements; however, their work is mostly empirical and focuses on lexical diversity. Further, the interdisciplinary focus of these papers is limited.
Yet, data diversity measurement has long been of central importance in a wide range of fields, including language learning and assessment (Jarvis, 2013), corpus linguistics (Kilgarriff, 2001; Egbert et al., 2022), sociology (Steele et al., 2022) and ecology (Simpson, 1949). Across these and other fields, scholars have also grappled with how to define and quantify diversity. Our position is that an interdisciplinary approach is essential for improving data diversity measurements in NLP.
@@
In this opinion paper, we provide a critical perspective on the fundamental challenge of measuring dataset diversity in NLP. We focus on language data and associated information (e.g., about the authors or annotators). We operationally define data diversity as the variability of the data across specific dimensions (e.g., semantic, syntactic, lexical, sociolinguistic, genre, language, etc.). Cross-lingual diversity has received increased attention in NLP Joshi et al. (2020); Ploeger et al. (2024), leading to standardized language diversity metrics (Ponti et al., 2020; Samardzic et al., 2024). Within-language diversity, however, remains underexplored. For this reason, we focus on measuring diversity in monolingual settings. We first motivate why data diversity should be measured (§2). Then, we structure our discussion around three core challenges: diversity dimensions (§3.1), their measurement (§3.2), and evaluation of measurement (§3.3). Throughout our paper we highlight connections with other fields. Rather than offering definitive answers, we hope to spark discussion and to encourage more rigorous and interdisciplinary approaches to measuring data diversity. We conclude with concrete recommendations for NLP research that uses or develops diversity measures (§4).
@@ 2 Why Measure Data Diversity?
Although many papers make qualitative judgements about the diversity of datasets (e.g., Zhou et al. 2023; Röttger et al. 2022), the exact nature and extent of the differences in diversity can remain unclear. We take the view that data diversity should be measured — that is, quantitatively characterized — along different dimensions (Lai et al., 2020; Mitchell et al., 2023; Zhao et al., 2024).
@@
First, data collection, especially when requiring human annotation, can be resource-intensive. Recent work (Quattoni and Carreras, 2021; Su et al., 2022; Shi et al., 2021; Alcoforado et al., 2024) shows that human annotation can be employed more effectively when selecting diverse datapoints. Second, including quantitative metrics on dataset diversity could further improve the documentation of datasets (McMillan-Major et al., 2024; Gebru et al., 2021) and contribute to dataset analysis tools (Lohr and Hahn, 2023; Ramponi et al., 2024). This could support early intervention, during data collection, instead of post hoc. For example, researchers could proactively collect more data from certain language varieties to improve the fairness of models. Third, leveraging data diversity measurements for training and prompting language models has led to better performing models, exhibiting improved robustness (Bukharin et al., 2024) and generalization (Levy et al., 2023). Lastly, when it comes to evaluation, insufficiently substantiated diversity claims have caused disagreement on the generalizability of results. For example, an evaluation dataset that covers “a diverse range of text genres” (Wang et al., 2018) may, by other standards, deemed “hardly diverse” (Raji et al., 2021). Systematic data diversity measurement could help researchers to assess and interpret the generalizability of evaluation results.
Measuring data diversity can open up new research directions
One recent example is understanding the effect of models being continuously trained on generated text through the lens of data diversity (Guo et al., 2024). Moreover, systematic diversity measurement holds potential for predicting a model’s task performance (Xia et al., 2020), which would allow researchers to better select datasets that align with their use cases, or predict potential fairness issues.
@@ 3 Challenges in Measuring Data Diversity
This section reflects on three key challenges when measuring data diversity: (i) What diversity dimensions should we consider? (§3.1); (ii) How should we measure them? (§3.2); (iii) How should we assess the quality of the measures? (§3.3).
@@ 3.1 Dimensions of Data Diversity
NLP studies have considered a range of diversity dimensions. Lexical diversity (e.g., Gandhi et al. 2024; Guo et al. 2024; Lohr and Hahn 2023; Shaib et al. 2025) and semantic diversity (e.g., Lai et al. 2020; Li et al. 2024; Yu et al. 2022) have featured in many studies. Research has also considered various other dimensions, including syntactic (Guo et al., 2024), topical (Schiller et al., 2024), and genre diversity (Liu and Zeldes, 2023). Furthermore, prior work has addressed task-specific dimensions, e.g., based on question types (Ghazaryan et al., 2025), key points in essays (Padmakumar and He, 2024) or summarization styles (Grusky et al., 2018).
@@
We cannot — and need not – measure all possible dimensions; likewise, datasets cannot (and need not) be highly diverse along all of them. Instead, we should selectively focus on the dimensions most relevant to the task at hand. With this broader, yet targeted approach to data diversity measurement, we can clarify which dimensions matter for which tasks, and make more informed decisions about dataset construction.
@@ 3.2 How Should We Measure Diversity?
Data diversity dimensions can be decomposed into different components. Ecology, with its long tradition of diversity measurement, often recognizes two core components of biodiversity: richness (i.e., the number of species) and evenness (i.e. how balanced a distribution of species is) (Magurran, 2004). Other components are sometimes also considered, like disparity (taking the similarity between species into account) (Daly et al., 2018). While ideas from ecology may not map perfectly onto NLP, especially for dimensions that are not easily captured by categorical data, it offers a useful lens for thinking about data diversity (Friedman and Dieng, 2023). For instance, in language learning, ideas from ecology have been applied to measuring lexical diversity (Jarvis, 2013). Similarly, for topical diversity, we could consider the number of distinct topics (richness), how the data is distributed over topics (evenness), and the semantic similarity between topics (disparity).
NLP researchers, however, are rarely explicit about which components of a diversity dimension are considered. This ambiguity can lead to inconsistent or misleading interpretations. For example, does increasing dataset size lead to greater diversity? If the diversity measure captures richness then adding more data generally helps—or at least doesn’t hurt. However, if a balanced distribution across categories (i.e. evenness) is targeted, simply scaling up the dataset may not improve diversity and could even reinforce imbalances.
@@ 3.3 Desiderata and Evaluation
What makes a good diversity measure? This question is rarely addressed, and diversity measures are rarely evaluated (with a few exceptions; Tevet and Berant 2021; Han et al. 2022; Yang et al. 2025). Yet without doing so, it is unclear whether a measure captures what is intended or improves on others.
@@
This is widely used in the social sciences to evaluate the quality of measurements for abstract concepts, like personality or self-esteem, where no clear ground truth is present (Trochim et al., 2016). Data diversity faces a similar challenge (Zhao et al., 2024). Two key indicators of measurement quality are (construct) validity (i.e., are we measuring what we intend to measure) and reliability (i.e., the consistency or repeatability of the measurements). Note that while measurement theory provides a useful toolbox, it was not designed for NLP, and translating it to NLP also poses challenges (Van der Wal et al., 2024).
To illustrate how measurement theory might be useful, consider assessing validity (Trochim et al., 2016). For example, content validity involves assessing whether the full scope of the concept is adequately covered. Suppose we aim to measure dialect diversity in a dataset: if the chosen measure covers only a small subset of major dialects, this could point to a problem with content validity. As another example, convergent validity can be assessed by comparing multiple measures that are intended to capture the same diversity dimension (Shaib et al., 2025). A final example is concurrent validity, which involves checking whether the measure can distinguish between groups known to differ; for example, we could compare datasets known to vary in their levels of diversity.
@@
The NLP community should also engage more deeply with the question of what makes a good diversity measure. Desiderata could be theoretical (Daly et al., 2018) but also practical like the interpretability or actionability of measures (Delobelle et al., 2024). For example, measures based on pairwise similarities between embeddings are increasingly popular. These measures are often transparent in their formulation, but their dependence on embedding models complicates interpretation; furthermore, while they indicate how similar instances are, they do not indicate what is missing from a dataset, making it more difficult to intervene. Lastly, the efficiency of a diversity metric may be an important desideratum. Shaib et al. (2025) point out that some metrics are prohibitively slow and therefore not suitable for large-scale application.
@@ 4 Conclusion
We have examined key challenges in measuring data diversity in NLP. We conclude with six recommendations: (i) Clearly define what is meant by data diversity in the context of a study; (ii) Make deliberate, transparent choices about how diversity is measured. They should be guided by the specific research goals and application; (iii) Develop measures that better capture the social dimensions of diversity; (iv) Critically examine what makes a good diversity measure. NLP needs clearer guidance on the desirable theoretical and practical properties of data diversity measures; (v) Evaluate the quality of data diversity measurement; (vi) Look beyond NLP. Many fields have developed approaches to measure diversity; we should learn from them.
@@
Diversity is a rich and complex concept. Although we take the view that researchers should measure the diversity in datasets, we acknowledge that any quantification necessarily results in a loss of nuance and information. In addition, not all dimensions of diversity are necessarily measurable. Engaging with qualitative perspectives on diversity is therefore important; it could bring a critical and context-sensitive lens to diversity in NLP datasets. We also believe that combining manual and automatic analyses of datasets is an important path forward (Prabhu and Birhane, 2020). Relatedly, we acknowledge that quantifying data diversity may lead to diversity metrics becoming objectives on their own. We caution against the blind optimization of diversity metrics, as the quality of a dataset depends on various factors. Purely optimizing for a diversity metric might overlook other important factors of dataset quality, such as the relevance of texts.
@@ Acknowledgments
We would like to thank the members of the NLP and Society Lab at Utrecht University for their feedback. This work is supported by the ERC Starting Grant DataDivers 101162980 and the Carlsberg Foundation under the Semper Ardens: Accelerate programme (CF21-0454). We have used ChatGPT to improve the grammar and phrasing of some sentences in this paper.

### v2

@@ 1 Introduction
Language models seem to exhibit remarkable natural language understanding and generation capabilities. Yet numerous studies have pointed towards serious flaws: they encode societal biases (Gallegos et al., 2024; Hofmann et al., 2024) and are sensitive to spurious correlations (Du et al., 2023). Many of these weaknesses have been traced back to the training data (Feng et al., 2023; McCoy et al., 2024). Moreover, as the size of these training datasets continues to increase, improvements in task performance show diminishing returns (Tirumala et al., 2023). Data size is often prioritized at the expense of data quality (Paullada et al., 2021). However, attention for the quality and composition of NLP datasets is growing (Albalak et al., 2024; Elazar et al., 2024; van Noord et al., 2025; Yu et al., 2024).
@@
Unfortunately, the concept of “data diversity” is challenging to define and measure, and researchers are rarely explicit about what they mean with data diversity (Zhao et al., 2024). But without clear definitions and measurements, it is difficult to make progress. Previous papers that advocate for measuring data diversity have left important aspects unaddressed. Mitchell et al. (2023) emphasize the importance of measuring data characteristics, though diversity in NLP data was not their main focus. Zhao et al. (2024) highlight the need to measure data diversity; however, their discussion mostly focuses on considerations during dataset construction, rather than diversity measurement itself. Shaib et al. (2025) call for standardizing text diversity measurements; however, their work is mostly empirical and focuses on lexical diversity. Estève et al. (2025) recently surveyed approaches to measure diversity in NLP; however, they focused less on data diversity, and they did not focus on the broader challenges that we outline. Further, the interdisciplinary focus of these papers is limited.
Yet, data diversity measurement has long been of central importance in a wide range of fields, including language learning and assessment (Jarvis, 2013), corpus linguistics (Kilgarriff, 2001; Egbert et al., 2022), sociology (Steele et al., 2022) and ecology (Simpson, 1949). Across these and other fields, scholars have also grappled with how to define and quantify diversity. Our position is that an interdisciplinary approach is essential for improving data diversity measurement in NLP.
@@
In this opinion paper, we provide a critical perspective on the fundamental challenge of measuring dataset diversity in NLP. We focus on language data and associated information (e.g., about the authors or annotators). We operationally define data diversity as the variability of the data across specific dimensions (e.g., semantic, syntactic, lexical, sociolinguistic, genre, language, etc.). Cross-lingual diversity has received increased attention in NLP Joshi et al. (2020); Ploeger et al. (2024), leading to standardized language diversity metrics (Ponti et al., 2020; Samardzic et al., 2024). However, since within-language diversity remains underexplored, we focus on measuring diversity in monolingual settings. Still, our recommendations are not specific to any language. We first motivate why data diversity should be measured (§2). Then, we structure our discussion around three core challenges: diversity dimensions (§3.1), their measurement (§3.2), and evaluation of measurement (§3.3). Throughout our paper we highlight connections with other fields. Rather than offering definitive answers, we hope to spark discussion and to encourage more rigorous and interdisciplinary approaches to measuring data diversity. We conclude with concrete recommendations for NLP research that uses or develops diversity measures (§4).
Related concepts
Related to data diversity is a dataset’s representativeness, which is widely discussed in corpus linguistics but still lacks a consensus definition (Egbert et al., 2022). In other fields, like statistics and fairness in machine learning, the term has also been used in many ways (Chasalow and Levy, 2021). One common view in corpus linguistics is that a representative corpus is a miniature of the target population’s language; some other views align more with how we define data diversity (Egbert et al., 2022; Stefanowitsch, 2020).
Another related concept is data bias, which refers to systematic distortions in the data (Olteanu et al., 2019). It often refers to societal bias, e.g., when the data reflects historical prejudices, or statistical bias, e.g., due to issues like sampling biases and measurement errors (Mitchell et al., 2021). Although data bias and data diversity are related, they are not the same. For example, reducing statistical bias can increase diversity (e.g., by including more social groups), and some efforts to increase data diversity (e.g., oversampling minority groups) could increase statistical bias.
@@ 2 Why Measure Data Diversity?
Although many papers make qualitative judgements about the diversity of datasets (e.g., Zhou et al. 2023; Röttger et al. 2022), the exact nature and extent of the differences in diversity can remain unclear. We take the view that data diversity should be measured — that is, quantitatively characterized — along different dimensions (Lai et al., 2020; Mitchell et al., 2023; Zhao et al., 2024). Measuring data diversity is relevant for both creating and reusing existing datasets, whether they are task-specific or more general-purpose.
@@
First, it would support more deliberate and informed data collection and curation (Rogers, 2021). Data collection, especially when requiring human annotation, can be resource-intensive. Recent work (Quattoni and Carreras, 2021; Su et al., 2023; Shi et al., 2021; Alcoforado et al., 2024) shows that human annotation can be employed more effectively when selecting diverse datapoints. Furthermore, quantitative metrics on dataset diversity could improve the documentation of new and existing datasets (McMillan-Major et al., 2024; Gebru et al., 2021) and contribute to dataset analysis tools (Lohr and Hahn, 2023; Ramponi et al., 2024). This could support early intervention, during data collection, instead of post hoc. For example, researchers could proactively collect more data from certain language varieties to improve the fairness of models. Second, leveraging data diversity measurements for training and prompting language models has led to better performing models, exhibiting improved robustness (Bukharin et al., 2024) and generalization (Levy et al., 2023). Lastly, when it comes to evaluation, insufficiently substantiated diversity claims have caused disagreement on the generalizability of results. For example, an evaluation dataset that covers “a diverse range of text genres” (Wang et al., 2018) may, by other standards, deemed “hardly diverse” (Raji et al., 2021). Systematic data diversity measurement could help researchers to assess and interpret the generalizability of evaluation results.
Measuring data diversity can clarify its impact and help explain and predict model behavior
Current knowledge of data diversity in NLP is fragmented and coarse-grained, partly because it is often not measured or measured inconsistently across studies. Rigorous data diversity measurement could reveal which dimensions of data diversity most strongly impact model behavior (e.g., task performance, generation quality, robustness, fairness), and how its effects interact with dataset size and other properties. For example, Guo et al. (2024) study the effect of models being continuously trained on generated text through the lens of data diversity. Besides explaining model behavior, diversity measurements could also serve as predictive signals. For example, they could help predict a model’s task performance (Xia et al., 2020), enabling researchers to better select datasets that align with their use cases, or anticipate fairness issues.
@@ 3 Challenges in Measuring Data Diversity
This section reflects on three key challenges when measuring data diversity: (i) What diversity dimensions should we consider? (§3.1); (ii) How should we measure them? (§3.2); and (iii) How should we assess the quality of the measures? (§3.3).
@@ 3.1 Dimensions of Data Diversity
NLP studies have considered a range of diversity dimensions. The most common are lexical diversity (Gandhi et al., 2024; Guo et al., 2024; Lohr and Hahn, 2023; Shaib et al., 2025) and semantic diversity (Lai et al., 2020; Li et al., 2024; Yu et al., 2022). Studies have also considered various other broad dimensions, including syntactic (Guo et al., 2024), topical (Schiller et al., 2024), and genre diversity (Liu and Zeldes, 2023), as well as more task-specific dimensions, including question types (Kim et al., 2023; Ghazaryan et al., 2025), key points in essays (Padmakumar and He, 2024) and summarization styles (Grusky et al., 2018).
@@
We cannot — and need not — measure all possible dimensions; likewise, datasets cannot (and need not) be highly diverse along all of them. Instead, we should selectively focus on the dimensions most relevant to the task at hand. With this broader, yet targeted approach to data diversity measurement, we can clarify which dimensions matter for which tasks, and make more informed decisions about dataset construction.
@@
A broader view: Consider the situational factors of texts
NLP could draw from corpus linguistics by measuring the distribution of situational factors (Staples et al., 2015), such as the mode of communication (spoken or written, synchronous or asynchronous). More broadly, diversity could be measured over registers, e.g., fiction, news articles, and academic writing (Goulart et al., 2020). However, in NLP datasets, such information is often not available, and would need to be additionally annotated or inferred (Egbert et al., 2015).
@@ 3.2 How Should We Measure Diversity?
Data diversity dimensions can be decomposed into different components. Ecology, with its long tradition of diversity measurement, often recognizes two core components of biodiversity: richness (i.e., the number of species) and evenness (i.e., how balanced a distribution of species is) (Magurran, 2004). Other components are sometimes also considered, like disparity (taking the similarity between species into account) (Daly et al., 2018). While ideas from ecology may not map perfectly onto NLP, especially for dimensions that are not easily captured by categorical data, it offers a useful lens for thinking about data diversity (Friedman and Dieng, 2023; Estève et al., 2025; Lion-Bouton et al., 2022). For instance, in language learning, ideas from ecology have been applied to measuring lexical diversity (Jarvis, 2013). Similarly, for topical diversity, we could consider the number of distinct topics (richness), how the data is distributed over topics (evenness), and the semantic similarity between topics (disparity).
NLP researchers, however, are rarely explicit about which components of a diversity dimension are considered. This ambiguity can lead to inconsistent or misleading interpretations. For example, does increasing dataset size lead to greater diversity? If the diversity measure captures richness then adding more data generally helps—or at least doesn’t hurt. However, if a balanced distribution across categories (i.e., evenness) is targeted, simply scaling up the dataset may not improve diversity and could even reinforce imbalances.
@@
A broader view: Beware of cross-lingual differences
Diversity measures often rely on specific tools or models (e.g., tokenizers, embedding models), and the availability and quality of such tools varies between languages (Blasi et al., 2022). Some measures cannot be applied to all languages. For example, measuring syntactic diversity using a dependency parser (Guo et al., 2024) is not possible if such a tool is lacking. Even when tools are available, the quality of diversity measurements can vary across languages. For example, Lion-Bouton et al. (2022) showed that errors in automatic lemmatization can artificially inflate diversity scores. Structural properties of languages can also influence the distribution of diversity scores. For instance, type-token ratio (TTR), often used to measure lexical diversity, was found to correlate with morphological complexity (Kettunen, 2014). Therefore, special care should be taken when applying and comparing diversity measures across languages.
@@ 3.3 Desiderata and Evaluation
What makes a good diversity measure? This question is rarely addressed, and diversity measures are rarely evaluated (with a few exceptions; Tevet and Berant 2021; Han et al. 2022; Yang et al. 2025; Zhang et al. 2025). Yet without doing so, it is unclear whether a measure captures what is intended or improves on others.
@@
In the social sciences, measurement theory is widely used to evaluate the quality of measurements for abstract concepts, like personality or self-esteem (Trochim et al., 2016). Its benefits have already been highlighted for measuring biases (Van der Wal et al., 2024) and evaluating AI systems (Wallach et al., 2025; Fang et al., 2025). It also holds potential for data diversity measurement (Zhao et al., 2024). Two key indicators of measurement quality are (construct) validity (i.e., are we measuring what we intend to measure) and reliability (i.e., the consistency or repeatability of the measurements). Although measurement theory provides a useful toolbox, translating its methods and concepts to NLP also poses challenges (Van der Wal et al., 2024).
To illustrate how measurement theory can inform the evaluation of a diversity measure’s validity, consider the following three complementary sources of construct validity evidence (Trochim et al., 2016). Content validity involves assessing whether the full scope of a concept is adequately covered. Suppose we aim to measure dialect diversity in a dataset: if the chosen measure covers only a small subset of major dialects, this could point to a content validity problem. Convergent validity is another example, which examines whether multiple measures for the same diversity dimension produce similar diversity scores (Shaib et al., 2025). Finally, criterion validity tests whether the measure behaves as expected according to an external benchmark—for example, by constructing datasets with known differences in diversity and verifying that the measure correctly ranks them. Together, these three validity indicators provide complementary evidence of a measure’s validity.
@@
The NLP community should also engage more deeply with the question of what makes a good diversity measure. Desiderata can be theoretical, for example that a measure should be continuous or clarity on the conditions under which it is maximal (Daly et al., 2018). Desiderata can also be practical like the interpretability or actionability of measures (Delobelle et al., 2024). For example, measures based on pairwise similarities between embeddings are increasingly popular. These measures are often transparent in their formulation, but their dependence on embedding models complicates interpretation; furthermore, while they indicate how similar instances are, they do not indicate what is missing from a dataset, making it more difficult to intervene. The efficiency of a diversity metric may also be an important practical desideratum. Shaib et al. (2025) point out that some metrics are prohibitively slow and therefore not suitable for large-scale application.
@@ 4 Conclusion
We have examined key challenges in measuring data diversity in NLP and conclude with six immediately actionable recommendations. These recommendations relate to what diversity aspects should be measured (§3.1), how they should be measured (§3.2), and the assessment of such measures (§3.3): (i) Clearly define what is meant by data diversity in the context of a study; (ii) Make deliberate, transparent choices about how diversity is measured. They should be guided by the specific research goals and application; (iii) Develop measures that better capture the social dimensions of diversity; (iv) Critically examine what makes a good diversity measure. NLP needs clearer guidance on the desirable theoretical and practical properties of data diversity measures; (v) Evaluate the quality of data diversity measurement; and finally (vi) Look beyond NLP. Many fields have developed approaches to measure diversity; we should learn from them.
@@
Diversity is a rich and complex concept. Although we take the view that researchers should measure a dataset’s diversity, we acknowledge that any quantification necessarily results in a loss of nuance and information. Moreover, not all dimensions of diversity are necessarily measurable. Engaging with qualitative perspectives on diversity is thus important; it could bring a critical and context-sensitive lens to diversity in NLP datasets. We also believe that combining manual and automatic analyses of datasets is an important path forward (Birhane and Prabhu, 2021). Relatedly, we acknowledge that quantifying data diversity may lead to diversity metrics becoming objectives on their own. We caution against the blind optimization of diversity metrics, as the quality of a dataset depends on various factors. Purely optimizing for a diversity metric might overlook other important factors of dataset quality, such as the relevance of texts.
@@ Acknowledgments
We would like to thank the members of the NLP and Society Lab at Utrecht University and the anonymous reviewers for their feedback. This work is supported by the ERC Starting Grant DataDivers 101162980 and the Carlsberg Foundation under the Semper Ardens: Accelerate programme (CF21-0454). We have used ChatGPT to improve the grammar and phrasing of some sentences in this paper.